OpenStudy (anonymous):

How much time does it take the bus going 18 mph in traffic to cover a 42.3 mile route? Round to the nearest tenth if necessary. 60.3 hours 0.4 hours 2.4 hours 24.3 hours

OpenStudy (lgbasallote):

like i said before \[speed = \frac{distance}{time}\] there fore... \[time = \frac{distance}{speed}\] \[time = \frac{42.3}{18}\] what is 42.3 divided by 18?

OpenStudy (jewlzme17):

it would be C. 2.4 hours